OpenStudy (anonymous):

Every day, there are 3 times more likes on an internet video of a cat which is modelled by the function c(n) = (3)n − 1, where n is the number of days since the video posted. On the first day, there were 143 likes. What is the function that shows the number of likes each day?

OpenStudy (anonymous):

c(n) = (3)(143)(n − 1) c(n) = (143)n − 1 c(n) = (3)143 − 1 c(n) = 143(3)n − 1